MODS2 Commissioning Night 0

29 December 2014 UT

Because of problems with SX and clouds, some DDT observations could not be done. An LBC-R observation was tried, but the target is close the moon and there are clouds. Closing the dome at 20 UT.

Earlier in the afternoon,I stopped and restarted the AGw and IE services. Tonight, before taking data, I stopped and restarted the calibans, but the red caliban did not come up; it did not find /dev/sde. Called Rick and we followed the procedures:
  • cntl-alt-del on M2.RC (using a button at the top of the java screen brought up by Raritan)
  • reboot mods2data --- it stalled on disk errors

Will look at replacing the hard drive on M2.RC on Monday.

-- %USERSIG{OlgaKuhn - 2014-12-29}%

Comments

 
Topic revision: r2 - 29 Dec 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback